A surveyor measures the distance across a straight river by the following method: Starting directly across from a tree on the opposite bank, he walks x = 116 m along the riverbank to establish a baseline. Then he sights across to the tree. The angle from his baseline to the tree is 𝜃 = 34.0°. How wide is the river?

Y = _______ m



1
Expert's answer
2022-01-06T10:49:02-0500

d=X=116m

Angle"(\theta)=34°"


We know that

"tan\theta=\frac{Y}{X}"

"Y=Xtan\theta"

Where X=d=116m

"Y=116\times tan 34°"

"Y=116\times0.6745=78.24m"